A control chart is to be established on a process producing refrigerators. The inspection unit is...

  1. A control chart is to be established on a process producing refrigerators. The inspection unit is one refrigerator, and a common chart for nonconformities is to be used. As preliminary data, 16 nonconformities were counted in inspecting 30 refrigerators.

    (a) What are the three-sigma control limits?
    (b) What is the α-risk for this control chart?
    (c) What is the β-risk if the average number of defects is actually 2 (i.e., if c = 2.0)? (d) Find the average run length if the average number of defects is actually 2.
Answer #1

Answer: Given that A data, with 16 non com formities were counted in inspecting 30 refrigerators. a) To find the three-sigma=0+{\-P(X=1) +P (X=2).. =;- $ 80.8833 (0:5333) 20.5333(0.5333)27 y!t2 = |-[0.3129+0:083401 = 1-0.3963 2.0.6037 c) To find ß-